### v2

# article.bib

[…]
article.bib
Perception-Guided EEG Analysis: A Deep Learning Approach Inspired by Level of Detail (LOD) Theory
BG Tong1 Corresponding author: dackmoon123@126.com
(1Department of Psychiatry, Inner Mongolia People’s Hospital, Hohhot, China)
1 Introduction
1.1 Research Background
In the rapidly evolving fields of neuroscience and artificial intelligence, the relationship between psychological phenomena and neural signals remains an unsolved mystery. Although our understanding of brain structure and neural activity has deepened, inferring individual psychological experiences from microscopic biological signals remains highly challenging. This involves both technical difficulties and philosophical paradoxes, such as whether psychological phenomena can be reduced to neural signals [luppi2021like]. This study proposes a new approach from EEG biofeedback to perceptual regulation, exploring the interaction between music rhythm and neural activity to understand and guide human psychological states.
In the field of EEG signal analysis and psychological phenomena research, achievements and challenges coexist [bashivan2015learning]. Traditional studies often use machine learning methods to identify EEG signal patterns to distinguish mental states or diseases, such as epilepsy detection. These methods extract time-domain and frequency-domain features and use traditional classifiers for detection, but their effectiveness in classifying complex psychological phenomena is limited [zhang2019deeplearningdecodingmental]. Although deep learning methods have brought new opportunities and improved classification accuracy in some studies, they still face limitations when dealing with psychological phenomena, such as ignoring the complexity and dynamics of perception and lacking consideration of individual differences [roy2019deeplearningbasedelectroencephalographyanalysis]. In the field of psychotherapy, EEG biofeedback technology has been explored, but there are shortcomings in personalized treatment [2019EEG].
In contrast, this study aims to follow a different path. Drawing on Gestalt psychology theory, it treats perception as the core unit of psychological phenomena, breaking through traditional reductionist thinking. By introducing LOD theory, the study proposes a new deep learning method to dynamically adjust the processing levels of EEG signals, improving the accuracy of perceptual state recognition and providing new pathways for personalized psychological therapy. At the same time, innovative data collection and processing methods are adopted to advance the development of this field.
1.2 Challenges in Modeling Psychological Phenomena
Traditional deep learning methods, such as Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs), have shown remarkable performance in pattern recognition and prediction tasks [bhatti2024comparativeanalysisdeeplearning]. However, they face three main limitations when applied to psychological phenomena: (1)Reductionism Issue: Psychological phenomena are diverse and dynamic, making it difficult to represent them using a single data stream or pattern; (2)Causality Assumption: Many models treat neural signals as direct causes of psychological phenomena, ignoring the possibility that knowledge may be a posteriori meaning structure; (3)Individual Differences: The same EEG signal topology may produce different subjective psychological experiences in different individuals.These limitations serve as the motivation for the design of this study, which will be elaborated in detail in the subsequent sections [10.3389/fncom.2024.1402689].
1.3 Inspiration and Reference from Gestalt Psychology
Gestalt psychology emphasizes the holistic nature of experience and behavior, asserting that the whole is greater than the sum of its parts. In visual perception, there are numerous organizational principles, such as the laws of proximity and similarity. In the field of deep learning for image recognition, this holistic perspective aligns to some extent with how neural networks process images by extracting features from the overall pixel distribution, similar to how the human visual system organizes visual information based on Gestalt principles [chen2024gestaltcomputationalmodelpersistent]. Additionally, the concept of "insight" in Gestalt psychology shares a similar underlying mechanism with the Transformer model’s ability to predict the next word in a sequence [10.3389/fcomp.2024.1419831].
Figure 1: Gestalt Perspective: Collapse of Brain Signal Structure into Specific Perception.
1.4 Recognition and Feedback of Biological Signals in Psychotherapy
In the field of psychotherapy, the recognition and feedback of EEG signals hold significant importance. EEG signals contain rich psychological information, with different mental states exhibiting unique EEG patterns. By collecting EEG signals and converting them into feedback, combined with big data and artificial intelligence technologies, it is possible to construct psychological models, predict psychological changes, and intervene, thereby opening new avenues for psychotherapy. This study draws on Gestalt psychology theory, positioning perception as the fundamental unit of psychological phenomena and the integrator of information. A single-channel EEG device is used to collect data, aiming to extract meaningful perceptual features from complex background noise and develop training methods to optimize the understanding of overall perceptual states.
1.5 Inspiration and Reference from LOD Theory
The Level of Detail (LOD) concept was initially developed for 3D modeling, where it optimizes resource allocation by dynamically adjusting the level of detail in rendered objects [luebke2002level]. This study introduces LOD theory into psychological modeling, treating EEG signals as multidimensional dynamic data and capturing core features related to perception through hierarchical processing [weng2024selfsupervisedlearningelectroencephalogramsystematic].
In LOD, the camera’s perspective determines the level of detail displayed, analogous to an individual’s subjective cognitive framework. Different people may focus on different aspects of the same neural signals, and the same person’s cognitive framework may vary across different moments or contexts. LOD reduces unnecessary complexity by optimizing the processing of critical information, similar to the "background-foreground separation" in perceptual processes. The brain filters sensory information to form a holistic perception, akin to the attention mechanism in deep learning. In deep learning models, LOD can be applied to dynamic hierarchical processing (low-level vs. high-level features), multi-perspective generation (similar to "multi-camera perspectives"), and simplification and generalization.
This study draws inspiration from LOD and applies it to the framework of deep learning algorithms, including the following operations:
1.5.1 1.Importance-Based LOD:
In techniques for optimizing 3D graphics rendering performance, importance-based LOD primarily considers the visual significance of objects in a scene. It not only focuses on the distance between the observer and the object but also takes into account factors such as the object’s size, color, texture complexity, and position in the frame. In this study, greater computational resources are allocated to significant EEG features to enhance the model’s sensitivity to key features. The data structure is designed as a two-dimensional combination of EEG data and audio data (with EEG data collected with a 200-millisecond delay). It is assumed that the important perceptual "points" depicted by LOD must coincide with the time points of audio-induced data, leveraging both the rhythm of audio nodes and the necessary "framing" of perception. The calculation formula is as follows:
$\hat{X}=X\cdot W_{i}$ (1)
where $Wi$ represents the learned feature importance weights.
1.5.2 2.Distance Threshold Method:
The distance threshold method determines the level of detail based on the distance between the observer (camera) and the rendered object. When the distance between the object and the camera exceeds a specific threshold, the system switches to a lower level of detail. In this study, the distance threshold method is applied as a time-series operation to simplify input signals and simulate perceptual features at different time scales. When collecting EEG frequencies, an average of 256 data points is collected per loop. As the loop rate varies, the sampling interval changes accordingly, allowing the density of EEG feature data to dynamically adjust with the rhythm density. This better "outlines" the perceptual picture. The calculation formula is as follows:
$\hat{X}=\text{Pool}(X,k)$ (2)
where $X$ is the input data,$Pool$ represents the pooling operation, and $k$ is the pooling window size.
1.5.3 3.Dynamic LOD:
Compared to methods using predefined distance thresholds, dynamic LOD technology allows for smoother adjustments to the level of detail of objects. It can continuously adjust the complexity of objects within a certain range, including vertex reduction and texture simplification. Dynamic LOD automatically adjusts the level of detail based on real-time conditions (such as changes in perspective, hardware performance, and scene complexity) to maintain optimal performance and visual quality. This study posits that perceptual experiences exhibit similar high-dimensional spatial dynamics in human consciousness. Therefore, reinforcement learning is introduced into the algorithm design. It is observed that while subjects believe they are immersed in a rhythm and maintain a certain perception during EEG data collection, their internal perspective actually undergoes changes. For example, the current perceptual experience is a perception of previous perceptions, and the EEG-depicted image resembles a "picture within a picture," echoing the level of detail and focal points influenced by object distance in dynamic LOD. By leveraging relaxation and focus data provided by the TGAM chip, deep learning is used for correlation analysis (score-based reinforcement unsupervised learning). The calculation formula is as follows:
$\hat{X}=f(X)\cdot\sigma(W_{e}\cdot E)$ (3)
where $E$ represents relaxation and focus levels, $W_{e}$ is the weight matrix,and $\sigma$ is the activation function.
1.5.4 4.Screen Space Error (SSE):
Screen Space Error is used to evaluate the geometric difference between a simplified 3D model and the original high-detail model, ensuring that the simplified model’s screen projection remains visually consistent with the original model and avoids noticeable discrepancies. This study hypothesizes that perceptual phenomena are essentially the reprojection of signal topologies generated by brain neurobiochemical processes (to be discussed in detail later). In the specific algorithm implementation, a CNN module is introduced, treating the subject’s perceptual experience during rhythm loops as "fragile fragments." The collected EEG feature values are transformed from time series into a 256×4×2 matrix and input into an adjusted CNN convolutional neural network. The calculation formula is as follows:
$\text{SSE}=\frac{1}{N}\sum_{i=1}^{N}(X_{i}-\hat{X}_{i})^{2}$ (4)
2 Methods
2.1 Data Collection and Preprocessing
2.1.1 Data Collection
This study focuses on collecting EEG data from the researchers themselves, aiming to explore individualized psychological perception and EEG characteristics from the perspective of Gestalt psychology. Given the uniqueness of individual psychological experiences, portable EEG devices were used for data collection to capture these unique EEG patterns and their underlying psychological structures.
The experiments were conducted in a quiet daily environment, with the researchers wearing TGAM chips for data collection. The TGAM chip plays a crucial role in this study’s data collection process. It not only records raw data in real time but also provides spectral information for $\alpha$, $\beta$, $\theta$, and $\delta$ waves, as well as two key indicators: relaxation and focus. These indicators range from 0 to 100, with higher values indicating greater relaxation or focus, providing an important basis for quantitatively assessing the psychological state of the subjects, see Table 1.
Byte Value Explanation
1 0xAA [SYNC]
2 0x20 [PLENGTH] (payload length) of 32 bytes
3 0x00 [POOR_SIGNAL] Quality
4 0x83 No poor signal detected (0/200)
8 0x18 [ASIC_EEG_POWER_INT]
9 0x20 [VLENGTH] 24 bytes
10 0x00 (1/3) Begin Delta bytes
11 0x00 (2/3)
12 0x94 (3/3) End Delta bytes
13 0x00 (1/3) Begin Theta bytes
14 0x00 (2/3)
15 0x42 (3/3) End Theta bytes
16 0x00 (1/3) Begin Low-alpha bytes
17 0x00 (2/3)
18 0x0B (3/3) End Low-alpha bytes
19 0x00 (1/3) Begin High-alpha bytes
20 0x00 (2/3)
21 0x64 (3/3) End High-alpha bytes
22 0x00 (1/3) Begin Low-beta bytes
23 0x00 (2/3)
24 0x4D (3/3) End Low-beta bytes
25 0x00 (1/3) Begin High-beta bytes
26 0x00 (2/3)
27 0x3D (3/3) End High-beta bytes
28 0x00 (1/3) Begin Low-gamma bytes
29 0x00 (2/3)
30 0x07 (3/3) End Low-gamma bytes
31 0x00 (1/3) Begin Mid-gamma bytes
32 0x00 (2/3)
33 0x05 (3/3) End Mid-gamma bytes
34 0x04 [ATTENTION] eSense
35 0x0D eSense Attention level of 13
36 0x05 [MEDITATION] eSense
37 0x3D eSense Meditation level of 61
38 0x34 [CHKsUM] (1’s comp inverse of 8-bit
Payload sum of 0xCB)
Table 1: The data structure in EEG packets sent by TGAM Bluetooth
Each data collection session lasted 6 to 8 minutes and included 200 rhythm loops. The rhythm speed could be adjusted between 90 and 140 BPM, with each adjustment made in increments of 5 BPM. For example, when the speed was set to 120 BPM, each rhythm loop lasted 2 seconds (60 seconds / 120 BPM = 0.5 seconds per beat, with 4 beats forming a loop, thus each loop lasting 2 seconds) [Yedukondalu2025]. Before each session, the researcher selected a seasonal label (spring, summer, autumn, winter) reflecting their current psychological state, which served as the training label for the deep learning model. A total of 418 independent data collection sessions were completed, accumulating a rich dataset of paired EEG and music rhythm data.
2.1.2 Data Preprocessing
First, the dataset was divided into training and test sets using an 80%-20% split. If necessary, a validation set could also be created. This division helps evaluate model performance and prevent overfitting. Each sample contains the following information:
  • EEG Data: Shape (N, 256, 4, 2), where N represents the number of samples.
  • Focus and Relaxation Scores: Shape (N, 2), recording the focus and relaxation scores for each sample.
  • Drum Parameters: Shape (N, 16), including drum speed and drum arrangement codes.
  • Perceptual Labels: Shape (N, 2), recording the state and its percentage (spring, summer, autumn, winter).
  • Operation Records: Shape (N, 1), recording adjustments to speed and density.
EEG data preprocessing is a critical step before model training. Raw EEG data was collected without frequency-domain transformation to preserve the maximum amount of information. The sampling frequency was set to 512 Hz, with 256 data points collected per rhythm loop, ensuring precise alignment with the temporal structure of the music rhythm.
For audio processing, music rhythm signals were synchronously collected, and the average audio value within the first 200 milliseconds of each rhythm loop was calculated as a feature input. This approach effectively captures the influence of rhythm on EEG signals while providing the model with contextual audio information.
Since the TGAM dry electrode chip was used, EEG signals may be affected by noise and artifacts, such as eye movement artifacts (blinking) and power line interference (50/60 Hz). To ensure signal quality, low-pass filtering and Independent Component Analysis (ICA) techniques were employed to remove these artifacts, enhancing signal purity, which is crucial for improving model accuracy.
Finally, EEG and audio data were integrated into a 256×4×2 data structure. This structure cleverly incorporates time-series information and corresponding audio features, serving as the input for the subsequent Convolutional Neural Network (CNN) model. This lays a solid foundation for the model to accurately analyze the relationship between EEG signals and music rhythm [pandey2021brain2depthlightweightcnnmodel].
2.2 Software Design and Development
The software system for this study was developed using the Unity engine, aiming to achieve interactive EEG biofeedback and music rhythm modulation. This design plays a crucial connecting role in the overall research methodology, tightly integrating data collection, model training, and user experience.
The software system incorporates several key elements. For audio materials and MIDI structure, MIDI drum materials generated by software such as BandInBox 2015, Addictive Drums 2, Studio Drummer, and EZdrummer 2 were used, totaling approximately 4GB. These materials cover a variety of music styles, including jazz, rock, and hip-hop, and exhibit a wide range of drum complexity, from simple to intricate. Each MIDI file is pre-labeled with important information such as BPM (beats per minute), drum density, and style tags. This enables the system to dynamically adjust music rhythms based on user feedback, providing a rich foundation for personalized music modulation based on user EEG data and psychological states.
The user interface and interaction design were implemented using Unity’s GUI system, featuring two adjustment axes: tension-relaxation and complexity-simplicity. Users can click buttons on the interface, such as "more tense," "more relaxed," "more complex," or "more simple," to adjust drum rhythms in real time. After each adjustment, the system automatically selects suitable drum combinations from the MIDI library and plays them using Unity’s audio engine, see Figure 2. This real-time interactive design not only enhances users’ ability to autonomously modulate music rhythms but, more importantly, creates diverse experimental conditions for collecting EEG data related to different psychological states through users’ active operations. This allows the research to more comprehensively explore the complex relationship between music rhythm and psychological states.
Figure 2: Schematic Diagram of the EEG Data Acquisition and Music Rhythm Modulation System
For the integration of EEG data with Unity, the TGAM chip establishes a stable connection with an Android device running the Unity program via Bluetooth, enabling real-time transmission of EEG data. Upon receiving the EEG data, the Unity program immediately synchronizes it with the current drum rhythm, generating a specific 256×4×2 data structure. This data structure serves as the input for the deep learning model, ensuring that the model can simultaneously analyze the relationship between EEG signals and music rhythms. This provides robust data support for accurately identifying and guiding perceptual states.
Additionally, Unity’s built-in reinforcement learning module (e.g., ML-Agents) was ingeniously applied to optimize drum rhythm adjustment strategies. The model dynamically adjusts drum parameters based on user EEG data (including key indicators such as relaxation and focus) and subjective feedback (e.g., clicking the "adapt to current state" button). Through continuous learning and optimization, the system gradually guides users toward the target perceptual state, achieving intelligent interaction between the software system and users. This provides an innovative data-driven approach to music rhythm modulation for the research.
2.3 Data Collection and Label Assignment
To ensure the validity of the data and the high quality of model training, we carefully designed a dynamic label assignment mechanism that runs through the entire data collection process. This mechanism is of great significance for accurately reflecting changes in users’ perceptual states.
2.3.1 Initial State Selection
At the beginning of each experiment, the subject is required to select a seasonal label ("spring," "summer," "autumn," or "winter") that reflects their current psychological state as the initial perceptual state. Data collection then continues for 200 rhythm loops (approximately 6–8 minutes). During the data collection process, the drum rhythm speed can be flexibly adjusted between 90 and 140 BPM, with each adjustment made in increments of 5 BPM. This ensures that EEG data is collected under different rhythm conditions. For example, when the drum rhythm speed is set to 120 BPM, each rhythm loop lasts 2 seconds. This standardized collection process facilitates subsequent data processing and analysis.
2.3.2 Initial State Selection
The core part of this mechanism is the state transition and label assignment. If the subject remains in the initially selected perceptual state at the end of the experiment, all collected data will be explicitly labeled as 100% initial state. For example, if the initial state is "autumn" and no state change occurs by the end of the experiment, all data will be labeled as 100% autumn. However, if the subject transitions from the initial state (e.g., "autumn") to another state (e.g., "summer") during the experiment, the data labels will gradually change from 100% initial state to 100% target state, forming a linear gradient process [Craik_2019]. This dynamic label assignment method precisely records the trajectory of the subject’s perceptual state changes during the experiment, providing the subsequent deep learning model with rich, time-series-labeled information. This helps the model better learn and understand the patterns of perceptual state transitions [ding2024eeg].
2.3.3 Data Storage and Usage
Each experiment generates 50 data structures of size 256×4×2. Each data structure contains important information such as EEG features, audio features, relaxation and focus scores, and perceptual state labels. These data comprehensively record various physiological and psychological indicators of the subject during the experiment and accurately reflect changes in perceptual states through the labels. These rich data are used to train and validate the deep learning model, ensuring that the model can accurately identify and guide perceptual states. This provides a solid data foundation for the entire study.
2.4 Deep Learning Model Design
The deep learning model designed in this study aims to extract features from EEG signals and music rhythms for the classification of perceptual states. The overall structure of the model consists of three main parts: data input and preprocessing, feature extraction and fusion, and perceptual state classification [10.3389/fnhum.2024.1442398].
2.4.1 Data Input and Preprocessing
  • The model input is a 256×4×2 tensor, representing the EEG and audio signals for each rhythm loop, where:
    • Time Dimension: 256 data points, ensuring alignment with the temporal structure of the music rhythm.
    • Loop Dimension: 4 rhythm loops form a complete cycle, capturing the periodic features of the rhythm.
    • Feature Dimension: 2, including EEG signals (Raw Data, $\alpha$, $\beta$, $\theta$, $\delta$ waves) and audio features (BPM, drum density).
  • Data Preprocessing Steps:
    • Normalization: Normalize EEG and audio features to unify data scales.
    • Smoothing: Use low-pass filtering to remove noise while retaining low-frequency EEG signals.
    • Data Augmentation: Introduce time-axis shifting and noise perturbation to enhance model robustness.
2.4.2 Feature Extraction and Fusion Based on LOD Theory
  • CNN Architecture: A multi-layer CNN is used to extract spatiotemporal features. The structure is as follows:
    • First Layer: 1×1 convolution to capture local features, reduce parameters, and output 32 channels with ReLU activation.
    • 3×3 convolution + max pooling to extract spatial features of time and audio signals, followed by ReLU activation and 2×2 max pooling.
    • 3×3 convolution + max pooling to further extract high-level features, capturing the complex relationship between EEG signals and music rhythms, followed by ReLU activation and 2×2 max pooling.
  • LOD Fusion:
    • Feature Fusion: After feature extraction by convolutional layers, EEG and audio features are concatenated along the channel dimension to form a unified feature representation.
    • LOD Application:
        Importance-Based LOD: Assign weights ($W_{i}$) based on feature importance to enhance the model’s sensitivity to key features. The formula is(1).
        Distance Threshold Method: Simplify signals through time-series pooling operations to simulate perceptual features at different time scales. The formula is(2).
        Dynamic LOD: Introduce reinforcement learning to dynamically adjust feature detail levels based on real-time conditions. The formula is(3).
2.4.3 Perceptual State Classification
  • Fully Connected Layer: Map the feature vectors extracted by the CNN to perceptual state classifications (spring, summer, autumn, winter). ReLU activation is used to introduce non-linearity.
  • Output Layer: Use the softmax function to generate probability distributions for each perceptual state. The output is the probability of each of the four states.
2.4.4 Model Training and Optimization
The model training process adopts a supervised learning strategy, with the following specific steps:
  • Loss Function:Use the cross-entropy loss function to optimize the classification task.The loss function is defined as:
    $L=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{C}y_{ij}\log(p_{ij})$ (5)
    where $L$ is the loss function, $N$ is the number of samples, $C$is the number of categories (4 in this study, corresponding to the four seasons), $y_{ij}$ is the true label of the $i$-th is the true label of the $p_{ij}$ is the model’s predicted probability of the $i$-th sample belonging to the $j$-th category.
  • Optimizer:
    • Use the Adam optimizer for parameter updates, with an initial learning rate of 0.001.
    • The Adam optimizer adaptively adjusts the learning rate, making it suitable for handling non-stationary data distributions.
  • Training Strategy:
    • Divide the dataset into training and test sets using an 80%-20% split to prevent overfitting.
    • Use early stopping to halt training when the validation loss no longer decreases.
2.5 Reinforcement Learning Framework
The reinforcement learning module dynamically adjusts music rhythms to guide users from their current perceptual state to the target state. This study employs a Deep Q-Network (DQN) combined with features extracted by the CNN and user feedback to optimize rhythm adjustment strategies [10220592], see Figure 3.
Figure 3: Flowchart of the DQN-based Perceptual State Guidance System.
2.5.1 State Space Definition
The state space represents the current environmental information of the system, including:
  • CNN-Extracted Features: Spatiotemporal features extracted from the 256×4×2 input data, representing the complex relationship between EEG signals and music rhythms.
  • Focus and Relaxation Scores: Focus and relaxation scores provided by the TGAM chip, ranging from 0 to 100, indicating the user’s current psychological state.
  • Drum Parameters: Current music rhythm parameters, including BPM (beats per minute), drum density, and style tags.
The state space is mathematically represented as:
$s_{t}=[f_{\text{CNN}},\text{Attention}_{t},\text{Relaxation}_{t},\text{BPM}_{t},\text{Density}_{t},\text{Style}_{t}]$ (6)
where $fCNN$ represents the feature vector extracted by the CNN, $\text{Attention}_{t}$ and $\text{Relaxation}_{t}$represent the focus and relaxation scores, and $\text{BPM}_{t}$,$\text{Density}_{t}$ and $\text{Style}_{t}$ represent the current BPM, drum density, and style tags, respectively.
2.5.2 Action Space Definition
The action space defines all possible operations that the model can perform. In this study, the action space includes the following:
  • Adjust BPM: Increase or decrease the speed of the music rhythm in steps of 5 BPM.
  • Adjust Drum Density: Increase or decrease the complexity of the drum pattern in steps of 1 level.
  • Switch Style: Switch between preset music styles (e.g., jazz, rock, hip-hop).
The action space is mathematically represented as:
[…]
2.5.3 Q-Learning Variant
This study employs a Deep Q-Network (DQN) as the reinforcement learning algorithm. Its core idea is to evaluate the expected return of each action through the Q-value function. The Q-value function is updated according to the Bellman equation:
$Q(s_{t},a_{t})=r_{t}+\gamma\max_{a_{t+1}}Q(s_{t+1},a_{t+1})$ (7)
where $r_{t}$ represents the immediate reward, $\gamma$ is the discount factor (typically set to 0.99), and $s_{t+1}$ and $a_{t+1}$ represent the next state and action, respectively.
The network structure of the DQN is as follows:
  • Input Layer: Receives the state space $s_{t}$ as input.
  • Hidden Layer: Consists of two fully connected layers, each with 256 neurons and ReLU activation.
  • Output Layer: Outputs the Q-value for each action, representing the expected return of taking a specific action in a given state.
2.5.4 Reward Function Design
To optimize the interaction between music rhythms and EEG signals for identifying and guiding individual perceptual states, this study designs a reward function based on user feedback and state transitions. The goal of the reward function is to encourage the system to guide subjects from their current state to the target state while penalizing state loss to prevent invalid data from affecting training.Reward Function Design Methods:
  • State-Matching Reward.The reward value is calculated based on the degree of match between the model’s predicted state and the subject’s confirmed state. The formula is:
    $R_{t}=$ $\lambda_{1}\cdot(1-\text{CrossEntropy}(p_{\text{model}},p_{\text{user}}))$ (8)
    $-\lambda_{2}\cdot\text{Loss}_{\text{transition}}$
    where $p_{\text{model}}$ represents the state distribution predicted by the model (e.g., 70% summer, 30% spring), $p_{\text{user}}~$ represents the state distribution confirmed by the user (e.g., 80% summer, 20% spring), $\text{Loss}_{\text{transition}}$ describes the loss function for whether the state transition trend aligns with the user’s choice, and $\lambda 1,\lambda 2$ are parameters controlling the reward weights.
  • State Transition Reward.When the state transitions from "autumn" to "summer," the reward is proportional to the change in label distribution. The formula is:
    $R_{t}=\alpha\cdot\Delta S_{t-1\rightarrow t}-\beta\cdot 1(\text{im})$ (9)
    where $\Delta S_{t-1\rightarrow t}$ represents the absolute change in label distribution (e.g., from [80% autumn, 20% summer] to [60% autumn, 40% summer]), $1(\text{im})$ is an indicator for state loss (1 if state loss occurs, otherwise 0), and $\alpha$,$\beta$ are weights for reward and penalty.
  • Rhythm Adjustment Effect Reward.Reward is given based on changes in relaxation and focus scores after the subject clicks the "adapt to current state" button. The formula is:
    $R_{t}=\gamma\cdot(\Delta\text{Relaxation}+\Delta\text{Concentration})$ (10)
    where $\Delta\text{Relaxation}$ and $\Delta\text{Concentration}$ represent the changes in relaxation and focus scores, respectively. If the scores improve, the reward is positive; if they decrease, a negative reward is given.
Through the design of the above reward functions, the model can dynamically adjust its strategy based on user feedback and state transitions, gradually optimizing the guidance effect.
2.5.5 Experience Replay Mechanism
To break the correlation between samples and improve training stability, this study implements an experience replay mechanism. The specific steps are as follows:
  • Experience Storage: Store each state transition ($s_{t},a_{t},r_{t},s_{t+1}$) in the replay buffer.
  • Random Sampling: Randomly sample mini-batches from the replay buffer for training, reducing the correlation between samples.
  • Target Network: Use a target network to calculate target Q-values, preventing fluctuations in Q-value estimation. The mathematical representation of experience replay is:
$L(\theta)=\mathbb{E}_{(s_{t},a_{t},r_{t},s_{t+1})\sim\mathcal{D}}\biggl{[}$ $\left(r_{t}+\gamma\max_{a_{t+1}}Q(s_{t+1},a_{t+1};\theta^{-})\right.$ (11)
$\left.-Q(s_{t},a_{t};\theta)\right)^{2}\biggr{]}$
where $\theta$ represents the parameters of the current network, $\theta^{-}$ represents the parameters of the target network, and $\mathcal{D}$ represents the replay buffer.
2.5.6 Training Strategy
The model training process follows these steps:
  • Initialization: Randomly initialize the DQN parameters and set the replay buffer size to 10,000.
  • Exploration and Exploitation: Use an $\epsilon$-greedy strategy to balance exploration and exploitation. The initial $\epsilon$ value is 1.0 and gradually decays to 0.01 during training.
  • Parameter Updates: Use the Adam optimizer to update network parameters with a learning rate of 0.001.
  • Target Network Updates: Copy the current network parameters to the target network at regular intervals (e.g., every 10,000 steps).
2.5.7 Guidance Mode Implementation
In guidance mode, the model dynamically adjusts music rhythms based on the learned strategy to guide users toward the target perceptual state. The specific steps are as follows:
  • State Recognition: The model identifies the user’s perceptual state based on current EEG data and music rhythms.
  • Action Selection: The optimal action (e.g., adjusting BPM, switching styles) is selected based on the Q-value function.
  • Real-Time Feedback: The user’s focus and relaxation scores, as well as subjective feedback (e.g., clicking the "adapt to current state" button), are used as the basis for adjusting the strategy.
3 Results
This section presents the experimental results of perceptual state recognition and guidance based on EEG signals. The experimental data is divided into two parts: first, evaluating the performance of the Convolutional Neural Network (CNN) in perceptual state classification tasks; second, evaluating the effectiveness of the Deep Reinforcement Learning (DQN) model in guiding perceptual states through music rhythms.
3.1 CNN Model Perceptual Classification Performance Evaluation
To evaluate the performance of the CNN model in EEG signal perceptual classification tasks, the collected EEG data was randomly divided into training and test sets using an 80%-20% split. The model was optimized on the training set and then evaluated on the test set. Evaluation metrics include accuracy, recall, F1 score, precision, as well as confusion matrix and ROC curve.
From Table 2, it can be seen that the CNN model achieved a high accuracy of 94.05% on the test set, indicating that the model can accurately classify EEG signals into corresponding perceptual states. Additionally, the recall and F1 score reached 94.05% and 0.940, respectively, demonstrating that the model achieved a balanced performance in detecting all perceptual categories. The precision of 94.32% indicates that among the samples predicted to belong to a certain perceptual state, a high proportion truly belong to that state.
Metric Value
Accuracy 94.05%
Recall 94.05%
F1-Score 0.940
Precision 94.32%
Table 2: shows the classification performance metrics of the CNN model on the test set.
Confusion Matrix
Figure 4: Confusion matrix of the CNN model on the test set.
The confusion matrix provides detailed information about the model’s classification performance across different categories. From the figure 4, it can be observed that the model performs well in distinguishing between "spring" and "summer" states, with only a small number of samples misclassified as "autumn". However, there is relatively more confusion between the "spring" and "winter" states, possibly due to the similarity in EEG features between these two seasons.
ROC Curve and AUC
Figure 5: ROC curves for each perceptual category and the average ROC curve of the CNN model.
ROC curves and AUC values are commonly used tools for evaluating model classification performance. From Figure 5, it can be observed that the AUC values for each perceptual state are close to 1, indicating strong classification capability. The average ROC curve has an AUC value of 0.94, further demonstrating the model’s excellent overall classification performance.
Conclusion
The above results show that the CNN model exhibits outstanding performance in EEG signal perceptual classification tasks, validating its ability to extract and classify perceptual-related features from EEG signals. The model can accurately classify different perceptual states with high accuracy, recall, F1 score, and AUC values.
3.2 DQN Model Perceptual Guidance Performance Evaluation
To evaluate the effectiveness of the DQN model in guiding perceptual states through music rhythms, we recorded the cumulative rewards and exploration rates during the training process and tested the model’s performance on the validation set.
Training Process
Figure 6: Cumulative Reward Curve of the DQN Model During Training
From Figure 6, it can be observed that as training progresses, the cumulative reward of the DQN model shows an upward trend and eventually stabilizes. This indicates that the model gradually learns to guide perceptual states by adjusting music rhythms.
Figure 7: Exploration Rate Decay Curve of the DQN Model During Training
Figure 7 shows that the model uses a high exploration rate in the early stages of training to try various actions. As training progresses, the exploration rate gradually decreases, indicating that the model begins to rely more on the learned strategies.
Validation Set Performance
On the validation set, the DQN model achieved a 92.45% success rate in guiding subjects’ EEG signals to the preset perceptual state. On average, it took 13.2 rhythm cycles (please replace with the actual average number of cycles) to complete a state guidance process, demonstrating the model’s good perceptual guidance capability and reasonable response speed.
Subjective Experience Evaluation
Although the EEG data shows that the DQN model achieved good results in perceptual guidance, subjective feedback from users revealed that approximately 50% of the researchers experienced psychological sensations corresponding to the target state during the rhythm adjustment process. This suggests that there is still room for improvement in the system’s effectiveness. Future enhancements could include introducing more personalized musical elements and more advanced reward functions.
Conclusion
The above results demonstrate that the DQN model, through reinforcement learning, can effectively adjust music rhythms to guide EEG activity, thereby influencing users’ perceptual states to some extent [chandanwala2024hybridquantumdeeplearning]. On the validation set, the model achieved a high success rate in guiding perceptual states. Although there is still room for improvement in subjective experiences, the results validate the potential of this method in EEG biofeedback and perceptual guidance.
4 Discussion
4.1 Experimental Limitations and Future Directions
During the design and implementation of this study, the conditions were relatively basic, and the process was somewhat rushed. Although some insightful results were obtained from the experimental data, there are still several limitations:
  • Single-Source Dataset: The dataset was limited to a single subject. Future work should expand to diverse subject groups to validate the model’s generalization ability.
  • Subjectivity of Labels: The subjective nature of the labels may limit classification performance. It is recommended to improve the labeling system by incorporating behavioral data or other physiological indicators.
  • Dynamic Adjustment of Reward Functions: The reward function in reinforcement learning needs to incorporate more dynamic adjustment mechanisms, such as real-time evaluation of user subjective feedback weights.
4.2 Evaluation of the Algorithm
The deep learning algorithm based on LOD theory proposed in this study demonstrates certain advantages in handling EEG analysis tasks for perceptual guidance. By introducing LOD theory into psychological modeling, we are able to approach the complex, multidimensional dynamic data of EEG signals from a new perspective. For example, in the model design, importance-based LOD technology enhances the model’s sensitivity to key EEG features by assigning higher weights to significant features, similar to focusing on important visual attributes in 3D modeling. The distance threshold method, applied to time-series pooling operations, simulates perceptual features at different time scales, allowing the density of EEG feature data to adjust dynamically with rhythm density, thereby better capturing the "outline" of perceptual scenes. Dynamic LOD technology, leveraging reinforcement learning, considers the dynamic nature of perceptual experiences in high-dimensional spaces and adjusts the level of detail based on real-time changes in the subject’s internal perspective, further optimizing model performance.
However, the algorithm also has some shortcomings. Current deep learning methods primarily focus on pattern matching at the sensory signal level. When dealing with psychological phenomena, due to a lack of deep understanding of the multi-level perceptual construction process, the predictive performance still has room for improvement. For instance, although we achieved some results in the experiments, there is still some confusion in distinguishing certain perceptual states (e.g., "autumn" and "winter"), possibly because we have not fully explored deeper perceptual information in EEG signals.
4.3 Psychological Phenomena Describable by the Algorithm
Perception, as the fundamental unit of psychological phenomena, has a unique multi-level structure. Sensation is the most basic neural signal, directly corresponding to data such as EEG, but it lacks higher-level psychological attributes. Perception, on the other hand, is the holistic interpretation of sensory signals, dependent on context and subjective experience [ma2024applying]. Emotions are composed of multiple perceptions, forming complex psychological states that can then be compressed into new perceptual units. Similarly, thinking is based on perception, and its results are also transformed into perceptual experiences. This multi-level relationship indicates that understanding the construction process of perception is crucial for accurately describing psychological phenomena.
4.4 Inspiration from Scaling Laws
In the era of large models, large language models rely on Scaling Laws (increasing data scale and model complexity). However, in the field of brain science, simply "stacking data and parameters" is not a viable strategy. The dimensionality of EEG data is not yet well-defined, and the intrinsic relationships within the data are holistic, making nonlinear patterns difficult to directly extract. This suggests that we need to find entry points to simplify complexity rather than blindly expanding data scale. LOD theory provides a perspective in this regard, capturing perceptual characteristics by "simplifying complexity" and extracting key patterns from massive data. This approach helps reduce reliance on large-scale annotated data while ensuring the robustness of perception, i.e., maintaining perceptual stability even when data is sparse or noisy [zhang2024investigatinggestaltprincipleclosure].
4.5 Philosophical Reflections
4.5.1 The Uncertainty Principle of Perception
Psychological phenomena (perception) are not solely determined by sensory input but are instead constructed through the interplay of the brain’s prior knowledge, subjective experience, and context. This characteristic is similar to the concept of superposition in quantum mechanics, where sensory data "collapses" into a specific psychological phenomenon only under a particular cognitive framework or perceptual perspective [rossi2024brain]. This reveals the subjectivity and unpredictability of human experience. Even if all neural signal data are available, without a "holistic framework," it is still impossible to reconstruct or predict individual psychological experiences, see Figure 1. Furthermore, our search for the causes of psychological phenomena may be misguided; psychological phenomena might be a priori existence, with neural activity serving as a post hoc rationalization.Although this perspective is somewhat paradigm-shifting, it provides a deeper direction for understanding psychological phenomena.
4.5.2 Collapse of Superposition States
Reinforcement learning can be seen as an implementation of the perceptual collapse mechanism. The system first generates multiple possible psychological perceptual outcomes (superposition states) from the input data. The user then determines a specific state through subjective feedback (e.g., selecting a label). The system uses this feedback to update its strategy, learning the user’s subjective perceptual patterns [keppler2024laying]. This reflects the time-dependent guiding nature of individual psychological phenomena. These findings suggest that we should adopt an "outcome-driven" learning approach in deep learning, using subjective perceptual labels to guide results while emphasizing the dynamic and unpredictable nature of psychological phenomena. This encourages models to focus on "holistic similarity" rather than "microscopic consistency."
4.5.3 Plato’s Allegory of the Cave
Plato’s Allegory of the Cave uses light projections as a metaphor for human cognition and perception, which aligns with this study’s reflections on psychological phenomena. Psychological phenomena are projections of the brain’s complex neural operations, serving as simplified and refined expressions that possess independence. Just as the shadows in the cave cannot fully represent the objects casting them, psychological phenomena cannot be fully reduced to neural mechanisms. In EEG research, we should respect the independence of psychological phenomena and use perceptual frameworks to construct a "cave wall"-like projection surface to interpret complex brain signals, rather than attempting to reconstruct all neural mechanisms.
4.5.4 Scientific Research and the Battle Against the Uncertainty Principle
Psychological phenomena may be projections of "absolute ideas," and scientific methods have limitations in understanding them, making it difficult to reach the "ideas" themselves. However, the significance of combining deep learning with brain science lies in finding the best projection mechanism to approximate psychological phenomena. This requires introducing philosophical unknowability into research design, accepting "uncertainty," and emphasizing simplification and abstraction.For example, the LOD model simplifies the complexity of brain signals, bringing the model closer to the projected form of psychological phenomena rather than overemphasizing microscopic details.
4.6 Future Directions
4.6.1 Future Improvements in Experimental Design
Although this study preliminarily validated the feasibility of guiding perceptual states by combining EEG biofeedback with music rhythm modulation, the experimental design still has some limitations. Specifically, in the selection of musical materials, the current drum rhythms, while effective in guiding users’ perceptual states, are limited by the simplicity of melodies and harmonies, which restricts the system’s expressive power. In future research, we plan to enrich the music library by introducing more diverse melodies and harmonic combinations to enhance the system’s guidance effectiveness [kim2024machine]. For example, by increasing the complexity of harmonies and the variability of melodies, the system can more precisely match users’ psychological states, thereby improving the user experience. Additionally, we aim to develop a more flexible user interface that allows users to customize music styles and rhythm parameters, further enhancing the system’s personalization and interactivity. We hope that these improvements will attract more researchers and users to explore the potential of EEG biofeedback and music modulation.
4.6.2 Incorporating Melodies and Harmonies in Future Studies
As artificial intelligence technology rapidly advances, brain-inspired computing is becoming a research hotspot [li2024brain]. Current AI algorithms, such as deep learning and reinforcement learning, have already simulated certain aspects of the human brain’s mechanisms, such as using neural networks to mimic neuronal activity. However, the development of AI is not merely a simple imitation of the human brain; it also drives a deeper understanding of how the brain works. For example, by studying the learning processes of AI models, we can better understand how the brain processes information, makes decisions, and forms perceptions. Conversely, as brain-inspired computing research progresses, we may discover new algorithms and models that further advance AI development. This bidirectional interaction makes brain-inspired computing and AI research a key focus for the future. We believe that, with technological advancements, brain-inspired computing will not only help us better understand the brain but also provide new ideas and methods for the development of AI [gigerenzer2024psychological].
5 Abbreviation
Full Name
AI Artificial Intelligence
CNN Convolutional Neural Network
DQN Deep Q-Network
EEG Electroencephalogram
GUI Graphical User Interface
ICA Independent Component Analysis
LOD Level of Detail
MLP Multilayer Perceptron
MIDI Musical Instrument Digital Interface
[…]